# Information-Theoretic Reward Decomposition for Generalizable RLHF

### v1

@@ 3.3 Guide Reward Learning with Prompt-free Reward
Figure 3: (Illustrative) We characterize training data samples in a 2-dimensional quadrant diagram, with the decomposed reward gaps $\Delta r_{1}$ and $\Delta r_{2}$. (a) shows initial data samples before training. (b) After training, the ideal distribution should be centered on the positive half of the $\Delta r_{1}$-axis, indicating that the preference depends solely on prompt-related information. (c-d) However, in the training process of $\Delta r_{\theta}$, the update following the BT model can only ensure the data points move in at least one of the positive directions of the $\Delta r_{1}$-axis or $\Delta r_{2}$-axis. If the update is based on samples with small $\Delta r_{2}$ (e.g. the left half), their movement upwards or to the right, along with other unupdated samples (e.g. the right half), will cause the distribution to be more centered on the positive $\Delta r_{1}$-axis. On the other hand, if the update is based on all samples, as shown in (d), the movement of samples with large $\Delta r_{2}$ values to the right exacerbates the existing prejudices, causing the distribution to become more centered on the positive $\Delta r_{2}$-axis.
@@
During the training process of $r_{\theta}$, one may consider leveraging the reward gap value $\Delta r_{\theta}(x,y_{1},y_{2})$ to selectively train more on samples with small reward gap (including negative ones). The reason is that, if the reward gap for $(x,y_{1},y_{2})$ satisfies $\Delta r_{\theta}(x,y_{1},y_{2})\approx 0$ or $<0$, it indicates that the current reward model $r_{\theta}$ struggles to distinguish between the preferred and dispreferred responses. Consequently, allocating more training budget to these samples can better align the reward model with human preference. Similar methods have been studied in training LLMs (Lin et al., 2024; Mindermann et al., 2022).
@@
Characterize ideal $r_{\theta}$ with $\Delta r_{1}$ and $\Delta r_{2}$ However, things become different if we consider prompt-related and prompt-free reward separately. Similar to the method discussed above, $\Delta r_{1}$ and $\Delta r_{2}$ are important properties of $r_{\theta}$ and can be leveraged to guide the training of $r_{\theta}$ itself. Specifically, during the learning process of $r_{\theta}$, we expect (i) $r_{\theta}$ to learn more prompt-related preference from the dataset (large prompt-related reward gaps); and (ii) $r_{\theta}$ to learn less prompt-free preference (small absolute values of prompt-free reward gaps). This is because both very large or very small prompt-free reward gaps (i.e., $\Delta r_{2}(x,y_{1},y_{2})\gg 0$ or $\ll 0$) indicate prejudice over the sampled responses, as they do not take the specific prompt into account. This prejudice will not only influence preferences under other prompts but may also introduce spurious reward gaps, e.g., reward gaps induced by properties only related to the responses, such as length bias (Dubois et al., 2024; Singhal et al., 2024). A more detailed discussion of this is given in §4.1.1.
@@
Prioritizing data with small $\Delta r_{2}$ To fulfill the expectations above, we present a straightforward method that prioritizes preference data in training reward models, with the aim of increasing prompt-related reward gaps while constraining the value of prompt-free reward gaps. Specifically, given data samples $\{(x,y_{1},y_{2})_{i}\}_{i=1}^{k}$ in each iteration, we first calculate their prompt-free reward gaps $\{\Delta r_{2}(x,y_{1},y_{2})_{i}\}_{i=1}^{k}$ based on Alg. 1, and then perform update for $r_{\theta}$ on those samples with small prompt-free reward gaps. For samples with large prompt-free gaps, we ignore them in this iteration and reinsert them into the buffer for subsequent updates.
@@
Analysis of prioritization mechanism.  Next, we provide an in-depth analysis for the proposed data prioritization mechanism. The standard update following the BT model only ensures the increase of $\Delta r_{1}(x,y_{w},y_{l})+\Delta r_{2}(x,y_{w},y_{l})$, so we cannot guarantee that either of them will necessarily increase. However, (i) if samples have small prompt-free reward gaps (e.g. $\Delta r_{2}<0$), regardless of whether their prompt-related reward gaps ($\Delta r_{1}$) are large or small, they should be used for updates. This is because an increase in $\Delta r_{1}$ suggests a better mastery of prompt-related preferences, while an increase in their $\Delta r_{2}$ indicates the elimination of existing prejudices, both of which are beneficial. (ii) Conversely, for samples with large $\Delta r_{2}$ (e.g., $\Delta r_{2}\gg 0$), if their $\Delta r_{1}$ are also large, updating is unnecessary since the prejudice led to gradient saturation in the BT model. If their $\Delta r_{1}$ are small, the preferences on them stem mainly from factors unrelated to the prompt (e.g., response length), which are easier for $r_{\theta}$ to learn from. Updating $r_{\theta}$ on them would exacerbate this issue, hindering generalization. Fig. 4 further illustrates this mechanism. Data samples with smaller $\Delta r_{2}$ help achieve a better trade-off between increasing prompt-related reward gaps and constraining prompt-free reward gaps.
@@
Practical Consideration. In practice, we design a binary-cluster mechanism to determine whether the value of $\Delta r_{2}$ is ‘small’ or ‘large’ in data prioritization. Specifically, at each step, we sample a batch of $(x,y_{1},y_{2})$ pairs from the preference dataset and perform one-dimensional binary clustering based on their $\Delta r_{2}$ values. Due to the dynamic nature of $r_{\theta}$ and the noise in data sampling, we do not directly use the obtained decision boundary for determination. Instead, we maintain a dynamic threshold, calculated as the exponential moving average of the boundary, and use this threshold for the determination. Due to limited space, we provide a detailed process in Appendix C.1 (Alg. 2).
@@ 4.1 Experiments on Manually Crafted Datasets
In this part, we manually construct datasets with specific characteristics from SHP (Ethayarajh et al., 2022), a commonly used open-source preference dataset. Based on these datasets, we compare the reward models trained with ordinary data and prioritized data. To thoroughly demonstrate their difference during training, we select four equally spaced training steps, sampling a number of data pairs and visualizing them in the same way in Fig. 4. Additionally, we evaluate the final reward model using Reward-Bench (Lambert et al., 2024) to assess its generalization capabilities. We choose LLaMA-3.2-1B-Instruct as the base model due to its lightweight nature and strong performance. For more details of the experiments, we refer to Appendix E.2.
@@ 4.1.1 Length-biased Dataset
The visualizations during training and the final results on Reward-Bench are shown in Fig. 4 and Table 2. If $r_{\theta}$ is trained with data uniformly sampled from $\mathcal{D}_{\rm bias}$ (Fig. 4, upper part), although its prompt-related reward gap ($\Delta r_{1}$) increases gradually, $r_{\theta}$ rapidly overfit to the length preference. This results in a significantly faster increase in $\Delta r_{2}$ for the chosen-longer pairs, and inevitably leads to a decrease in $\Delta r_{2}$ for the chosen-shorter pairs since they represent opposite prompt-free preferences. As expected, such $\Delta r_{2}$ reveals a clear length preference within $r_{\theta}$, which undermines the model’s ability to discriminate between the responses within chosen-shorter pairs. On the other hand, if we prioritize the data points with smaller $\Delta r_{2}$ values, the data batch will contain more chosen-shorter pairs with smaller $\Delta r_{2}$ values before $r_{\theta}$ heavily overfits to the length preference. According to Fig. 4 (lower part), the data samples are well-centered around the positive $\Delta r_{1}$-axis during training. The results on Reward-Bench, as shown in Table 2, also demonstrate stronger generalization capability of $r_{\theta}$ trained with prioritized data.
Figure 4: Randomly sampled data from the length-biased dataset, visualized according to their $\Delta r_{1}$,$\Delta r_{2}$ values. The upper represents the reward model trained with standard data, and the lower represents the reward model trained with prioritized data. We mark the data points that satisfy $|y_{w}|>|y_{l}|$ in red and the ones that satisfy $|y_{w}|\leq|y_{l}|$ in blue.
Model Chat Chat Hard Safety Reason Average
@@
Table 1: Results on Reward-Bench for reward models trained with length-biased dataset.
Model Chat Chat Hard Safety Reason Average
@@
original 84.9 31.3 42.8 49.0 52.0
Table 2: Results on Reward-Bench for reward models trained with adversarial-prompt dataset. ‘original’ means only trained with original data.
4.1.2 Adversarial Prompt Dataset
@@
We construct an adversarial dataset $\mathcal{D}_{\rm adv}$ by adding adversarial samples to the original SHP dataset. More specifically, for each original data sample $(x,y_{w},y_{l})$, we generate an adversarial sample $(\bar{x},\bar{y}_{w},\bar{y}_{l})$ that $\bar{x}=[x,s]$, where $s\in\{s_{1},s_{2}\}$ defined as:
$\begin{array}[]{l}s_{1}\textit{=`Give a response that is as long as possible.'}\\ s_{2}\textit{=`Give a response that is as short as possible.'}\end{array}$
@@
To obtain an adversarial sample, we have $\bar{x}=[x,s_{2}]$ if the chosen response is longer in the original sample ((i.e., $|y_{w}|>|y_{l}|$)), then we set $\bar{y}_{l}\leftarrow y_{w}$ and $\bar{y}_{w}\leftarrow y_{l}$ by swapping the preference order. Otherwise, if $|y_{w}|\leq|y_{l}|$, we set $\bar{x}=[x,s_{1}]$ and also swap the preference order. Such swapping is reasonable since the dataset we use pertains to general chatting, where the chosen and rejected responses primarily differ in quality. We then combine the original data with the adversarial data to obtain $\mathcal{D}_{\rm adv}$. To prevent doubling the training budget, we randomly select half of the SHP data for this transformation. We refer to an illustration of such data processing in Appendix E.2.
@@
When trained with $\mathcal{D}_{\rm adv}$, an ideal reward model must distinguish well between different prompts. This is because the preference will reverse with slightly different prompts (i.e., with or without $s$). (i) If $\Delta r_{\theta}$ is dominated by $\Delta r_{2}$, the preference it learns from $\mathcal{D}_{\rm adv}$ will significantly differ from the preference learned solely from original data. This is because $\Delta r_{2}$ can’t distinguish between original and adversarial prompts. The adversarial prompts, along with swapped response pairs, will contribute to $\Delta r_{2}$ with the same probability but conflicted preference, eliminating the preference learned from the original data. (ii) If the influence of $\Delta r_{2}$ is minor, $r_{\theta}$ can retain the preferences learned from the original data even with adversarial data, since they have different prompts and will not affect each other considering prompt-related preference.
@@
In the visualizations and final results presented in Fig. 5 and Table 2, it is evident that when $r_{\theta}$ is directly trained on $\mathcal{D}_{\rm adv}$, $\Delta r_{\theta}$ is dominated by $\Delta r_{2}$. Moreover, $r_{2}$ exhibits a strong preference for the originally rejected responses. This occurs because the simplified preference in the adversarial dataset can be inferred solely from response length, which is easier for the model to overfit. The preference of the overfitted reward closely mirrors the adversarial data. In contrast, our method prioritizes original data with more complicated preference, which is harder to learn during the reward training. This makes both the original and adversarial samples concentrated along the positive $\Delta r_{1}$-axis and results in a $r_{\theta}$ that demonstrates strong generalization capabilities.
Figure 5: Randomly sampled data during the training process under adversarial prompt dataset, the upper represents $r_{\theta}$ trained with standard data, the lower represents $r_{\theta}$ trained with prioritized data. We mark adversarial data in red and original data in blue.
@@ 4.2 Standard Experiments on Open-source Datasets
In this section, we evaluate our method using several commonly used open-source preference datasets. Unlike the manually crafted datasets discussed in §4.1, the prompt-free preferences of $r_{\theta}$ on these datasets may not directly exhibit specific features. However, a large prompt-free reward gap often suggests that $r_{\theta}$ may overfit to preferences that are irrelevant to the prompt. Such prompt-free preferences can negatively impact the generalization capabilities of $r_{\theta}$. We validate the effectiveness of our method across different base models and datasets through both direct evaluation of reward model accuracy and assessment of the induced policy performance. For more details of the experiments, we refer to Appendix E.2.
Evaluate Reward Model Accuracy. We use Reward-Bench (Lambert et al., 2024) as the benchmark to evaluate the reward model in a direct way, by testing its accuracy on well-organized but out-of-distribution queries. For a detailed illustration, we conduct experiments using both LLaMA-3-8B-Instruct (Dubey et al., 2024) and Mistral-7B-Instruct (Jiang et al., 2023) as the backbone of the reward model. For the training data, we use the SHP dataset, which is commonly used in RLHF.
The results are listed in Table 3. We choose vanilla BT reward training and RRM (Liu et al., 2024) as the baseline methods. RRM studies generalizable reward learning via causal inference and proposes a solution also from the data distribution perspective. Compared with baseline methods, our method demonstrates consistent improvements over different aspects of Reward-Bench. In Appendix E.3, we provide explanations of the superiority of the reward model trained by our method, compared with vanilla model and reward model trained with RRM.

### v2

@@ 3.3 Guide Reward Learning with Prompt-free Reward
During the training process of $r_{\theta}$, one may consider leveraging the reward gap value $\Delta r_{\theta}(x,y_{1},y_{2})$ to selectively train more on samples with small reward gap (including negative ones). The reason is that, if the reward gap for $(x,y_{1},y_{2})$ satisfies $\Delta r_{\theta}(x,y_{1},y_{2})\approx 0$ or <0<0, it indicates that the current reward model $r_{\theta}$ struggles to distinguish between the preferred and dispreferred responses. Consequently, allocating more training budget to these samples can better align the reward model with human preference. Similar methods have been studied in training LLMs [22, 26].
@@
Characterize ideal $r_{\theta}$ with $\Delta r_{1}$ and $\Delta r_{2}$ However, things become different if we consider prompt-related and prompt-free reward separately. During the learning process of $r_{\theta}$, we expect (i) $r_{\theta}$ to learn more prompt-related preference from the dataset (large $\Delta r_{1}(x,y_{1},y_{2})$); and (ii) $r_{\theta}$ to learn less prompt-free preference (small absolute values of $\Delta r_{2}(x,y_{1},y_{2})$). This is because when $\Delta r_{2}(x,y_{1},y_{2})\gg 0$ or $\ll 0$, it indicates strong prejudice over the sampled responses, as $\Delta r_{2}$ doesn’t take the specific prompt into account. This prejudice may introduce spurious reward gaps (e.g., reward gap induced by length bias)[13, 35].
@@
Prioritizing data with small $\Delta r_{2}$ Similar to the method discussed at the beginning of this section, $\Delta r_{1}$ and $\Delta r_{2}$ are important properties of $r_{\theta}$ and can also be leveraged to guide the training of $r_{\theta}$. To fulfill the desiderata above, we present a straightforward method that prioritizes preference data when training reward models, with the aim of increasing prompt-related reward gaps while constraining the value of prompt-free reward gaps. Specifically, given data samples $\{(x,y_{1},y_{2})_{i}\}_{i=1}^{k}$ in each iteration, we first decompose $r_{\theta}$ to get their prompt-free reward gaps $\{\Delta r_{2}(x,y_{1},y_{2})_{i}\}_{i=1}^{k}$ based on Alg. 1, and then perform update for $r_{\theta}$ on those samples with small prompt-free reward gaps. For samples with large prompt-free gaps, we ignore them in this iteration and reinsert them into the buffer for subsequent updates.
@@
Analysis of prioritization mechanism.  Next, we provide an in-depth analysis for the proposed data prioritization mechanism. The standard update following the BT model only ensures the increase of $\Delta r_{1}(x,y_{w},y_{l})+\Delta r_{2}(x,y_{w},y_{l})$, so we cannot guarantee that either of them will necessarily increase. However, (i) if samples have small prompt-free reward gaps (e.g. $\Delta r_{2}<0$), regardless of whether their prompt-related reward gaps ($\Delta r_{1}$) are large or small, they should be used for updates. This is because an increase in $\Delta r_{1}$ suggests a better mastery of prompt-related preferences, while an increase in their $\Delta r_{2}$ indicates the elimination of existing prejudices, both of which are beneficial. (ii) Conversely, for samples with large $\Delta r_{2}$ (e.g., $\Delta r_{2}\gg 0$), if their $\Delta r_{1}$ are also large, updating is unnecessary since the prejudice led to gradient saturation in the BT model. If their $\Delta r_{1}$ values are small, their $\Delta r_{\theta}$ values are dominated by factors unrelated to the prompt (e.g., response length). Updating $r_{\theta}$ on them would exacerbate this issue, hindering generalization. Fig. 3 further illustrates this mechanism. Data samples with smaller $\Delta r_{2}$ help achieve a better trade-off between increasing prompt-related reward gaps and constraining prompt-free reward gaps.
Figure 3: (Illustrative) We characterize training data samples in a 2-dimensional quadrant diagram, with the decomposed reward gaps $\Delta r_{1}$ and $\Delta r_{2}$. (a) shows initial data samples before training. (b) After training, the ideal distribution should be centered on the positive half of the $\Delta r_{1}$-axis, indicating that the preference depends solely on prompt-related information. (c-d) However, in the training process of $\Delta r_{\theta}$, the update following the BT model can only ensure the data points move in at least one of the positive directions of the $\Delta r_{1}$-axis or $\Delta r_{2}$-axis (up or right). If the update is based on samples with small $\Delta r_{2}$ (e.g. the left half), their movement upwards or to the right, along with other unupdated samples (e.g. the right half), will cause the distribution to be more centered on the positive $\Delta r_{1}$-axis. On the other hand, if the update is based on all samples, as shown in (d), the movement of samples with large $\Delta r_{2}$ values to the right exacerbates the existing prejudices, causing the distribution to become more centered on the positive $\Delta r_{2}$-axis.
@@
Practical Consideration. In practice, we design a binary-cluster mechanism to determine whether the value of $\Delta r_{2}$ is ‘small’ or ‘large’ in data prioritization. Specifically, at each step, we sample a batch of $(x,y_{1},y_{2})$ pairs from the preference dataset and perform one-dimensional binary clustering based on their $\Delta r_{2}$ values. Due to the dynamic nature of $r_{\theta}$ and the noise in data sampling, we do not directly use the obtained decision boundary for determination. Instead, we maintain a dynamic threshold, calculated as the exponential moving average of the boundary, and use this threshold for the determination. Due to limited space, we provide a detailed process in Appendix C.1 (Alg. 2). We justify the use of binary clustering and the EMA threshold in Appendix E.3 and demonstrate their robustness through experiments.
@@ 4.1 Experiments on Manually Crafted Datasets
In this part, we manually construct datasets with specific characteristics from SHP [14], a commonly used open-source preference dataset. Based on these datasets, we compare the reward models trained with ordinary data and prioritized data. To thoroughly demonstrate their difference during training, we select four equally spaced training steps, sample a number of data pairs at random, and visualize them by their $\Delta r_{1},\Delta r_{2}$ values using the same method as in Fig. 3. Additionally, we evaluate the final reward model using Reward-Bench [19] to assess its generalization capabilities. We choose LLaMA-3.2-1B-Instruct as the backbone due to its lightweight nature and strong performance. For more details of the experiments, we refer to Appendix E.4.
@@ 4.1.1 Length-biased Dataset
The visualizations during training and the final results on Reward-Bench are shown in Fig. 4 (a) and Table 1 (a). If $r_{\theta}$ is trained with data uniformly sampled from $\mathcal{D}_{\rm bias}$ (Fig. 4 (a), top), although its prompt-related reward gap ($\Delta r_{1}$) increases gradually, $r_{\theta}$ rapidly overfit to the length preference. This results in a significantly faster increase in $\Delta r_{2}$ for the chosen-longer pairs, and inevitably leads to a decrease in $\Delta r_{2}$ for the chosen-shorter pairs since they represent opposite prompt-free preferences. As expected, such $\Delta r_{2}$ indicates a clear length preference within $r_{\theta}$, which undermines $r_{\theta}$’s generalization capability. On the other hand, if we prioritize the data samples with smaller $\Delta r_{2}$ values, the training data batch will contain more chosen-shorter pairs with smaller $\Delta r_{2}$ values. According to Fig. 4 (a) bottom, the data samples are well-centered around the positive $\Delta r_{1}$-axis during training. The results on Reward-Bench, as shown in Table 1 (a), also demonstrate stronger generalization capability of $r_{\theta}$ trained with prioritized data.
Figure 4: (a): Visualizations for the length-biased dataset. We mark the data points that satisfy $|y_{w}|>|y_{l}|$ in red and the ones that satisfy $|y_{w}|\leq|y_{l}|$ in blue. (b): Visualizations for the adversarial prompt dataset. We mark adversarial data in red and original data in blue. In both (a) and (b), the top shows ordinary training data, and the bottom shows prioritized training data.
Model Chat Chat Hard Safety Reason Avg
@@
Model Chat Chat Hard Safety Reason Avg
@@
Table 1: (a): Results on Reward-Bench with length-biased data. (b): Results on Reward-Bench with adversarial-prompt data. “original” means vanilla reward training with only original data. Due to limited space, we present the mean results of 3 runs with different seeds. See Appendix E.1 for full results with std values.
4.1.2 Adversarial-prompt Dataset
@@
We construct an adversarial-prompt dataset $\mathcal{D}_{\rm adv}$ by adding adversarial samples to the original SHP dataset. More specifically, for each original data sample $(x,y_{w},y_{l})$, we generate an adversarial sample $(\bar{x},y_{l},y_{w})$ and add it to the dataset. $\bar{x}=[x,s]$ is the concatenation of the original prompt string $x$ and another string $s\in\{s_{1},s_{2}\}$. $s_{1},s_{2}$ are defined as:
[…]
@@
To ensure reasonable preference in the adversarial sample, we have $\bar{x}=[x,s_{2}]$ if the chosen response is longer in the original sample ((i.e., $|y_{w}|>|y_{l}|$)), and set $\bar{x}=[x,s_{1}]$ if if $|y_{w}|\leq|y_{l}|$. To prevent doubling the training budget, we randomly select half of the SHP data for this transformation. We refer to an illustration of such data processing in Appendix E.4.
@@
When trained with $\mathcal{D}_{\rm adv}$, an ideal $\Delta r_{\theta}$ should not be dominated by its prompt-free part ($\Delta r_{2}$). This is because the preference label in the dataset will reverse with slightly different prompts (i.e., with or without $s$). (i) If $\Delta r_{\theta}$ is dominated by $\Delta r_{2}$, the preference it learns from $\mathcal{D}_{\rm adv}$ will significantly degenerate compared with the preference learned solely from original data. This is because $\Delta r_{2}$ learns from both original and adversarial samples and overlooks their difference in the prompts. The adversarial samples will contribute to $\Delta r_{2}$ with the same probability but conflicted preference, eliminating the preference learned from the original data. (ii) If the influence of $\Delta r_{2}$ is minor, $r_{\theta}$ can retain the preferences learned from the original data even with adversarial data, since they have different prompts and will not affect each other considering prompt-related preference.
@@
In the visualizations and final results presented in Fig. 4 (b) and Table 1 (b), it is evident that when $r_{\theta}$ is trained on $\mathcal{D}_{\rm adv}$ with ordinary data, $\Delta r_{\theta}$ is dominated by $\Delta r_{2}$. Moreover, $r_{2}$ exhibits a strong preference for the originally rejected responses. This occurs because the simplified preference in the adversarial dataset can be inferred solely from response length, which is easier for the model to overfit. The preference of the overfitted reward closely mirrors the adversarial data. In contrast, our method prioritizes original data, which is harder to learn compared with adversarial samples. This makes both the original and adversarial samples concentrated along the positive $\Delta r_{1}$-axis and results in a more generalizable $r_{\theta}$. The performance of such $r_{\theta}$ is closer to that trained on the original data.
@@ 4.2 Standard Experiments on Open-source Datasets
In this section, we evaluate our method using several commonly used open-source preference datasets. Unlike the manually crafted datasets discussed in section 4.1, the prompt-free preferences of $r_{\theta}$ on these datasets may not directly exhibit specific features. However, a large prompt-free reward gap often suggests that $r_{\theta}$ may overfit to preferences that are irrelevant to the prompt. Such prompt-free preferences can negatively impact the generalization capabilities of $r_{\theta}$. We validate the effectiveness of our method across different base models and datasets through both direct evaluation of reward model accuracy and assessment of the induced policy performance. For more details of the experiments, we refer to Appendix E.4.
Evaluate Reward Model Accuracy. We use Reward-Bench [19] as the benchmark to evaluate the reward model in a direct way, by testing its accuracy on well-organized but out-of-distribution queries. For a detailed illustration, we conduct experiments using both LLaMA-3-8B-Instruct [12] and Mistral-7B-Instruct [18] as the backbone of the reward model. For the training data, we use a randomly sampled subset of 300K examples from the RLHFlow preference dataset [^2] https://huggingface.co/datasets/RLHFlow/pair_preference_model_dataset (originally 700K). We also conduct the same experiments on the SHP dataset (see Appendix E.2 for results).
The results are listed in Table 2 (a). We choose vanilla BT reward training and RRM [23] as the baseline methods. RRM studies generalizable reward learning via causal inference and proposes a solution also from the data distribution perspective. Compared with baseline methods, our method demonstrates consistent improvements over different aspects of Reward-Bench. In Appendix E.5, we provide explanations of the superiority of our method compared with baseline methods.
(a) Reward-Bench